# Membrane phononic integrated circuits

### v1

@@ IX.2 Example: Intracavity amplitude enhancement
Consider a square resonator of width $W_{\mathrm{res}}$ bookended by two waveguides of width $W_{\mathrm{wg}}$, as illustrated in Fig. 28. The resonator is impedance matched so that $\gamma_{1}=\gamma_{2}=\gamma/2$, were $\gamma_{i}$ is the coupling rate of the resonator to the $i^{\mathrm{th}}$ waveguide and $\gamma$ is the total decay rate of the resonator (we ignore other, intrinsic losses, i.e. $\gamma_{0}\approx 0$).
Consider the case where there is a drive in one waveguide ($a_{\mathrm{in,1}}=a_{\mathrm{in,2}}=0$), and unidirectional propagation. The drive frequency $\Omega$ is in the single mode regime of the waveguides and equals the resonator fundamental eigenfrequency ($\Delta=0$). Plugging the assumptions into Eq. (97) we find:
@@
We want to know the amplitude enhancement inside the resonator, that is $Z_{0,\mathrm{res}}/Z_{\mathrm{0,\mathrm{wg}}}$. We can substitute analytic expressions for the energy in the resonator and power in the waveguide into the normalisation conditions for $a$ and $a_{\mathrm{in}}$:
@@
Here $v_{g}$ is the group velocity in the waveguide.
@@ IX.3 Example: Filtering
[…]
@@ X Numerically finding coupling rates
Numerical methods are needed because because phononic membranes can be fabricated into a large variety of shapes that resist analytic treatment. As described in Section 9, the acoustic confinement provided by the impedance mismatch between the suspended membrane and the bulk substrate is very large (from the combination of contrasting speeds of sound and from the large difference in mass between the membrane and substrate [14]) enabling low-loss acoustic wave transmission as experimentally demonstrated in our earlier work with silicon nitride membranes [12]. This low loss—much lower than what is achievable for conventional semiconductor photonic high-index waveguides [135] or phononic high-index contrast circuits[136, 5]—means that sharp bends and abrupt changes in geometry can be realised while still achieving extremely high quality factors. Indeed, a typical refractive index mismatch in silicon photonics is $n_{\mathrm{Si}}-n_{\mathrm{vacuum}}\simeq 3.5-1=2.5$, corresponding to an impedance mismatch (for a plane electromagnetic wave in a nonconductive, nonmagnetic material) of $1/2.5=0.4$. In contrast, we saw in Section I.2 that silicon nitride membranes can have more severe impedance mismatches with the substrate of $Z_{\mathrm{mem}}/Z_{\mathrm{sub}}\simeq 0.06$, even before considering thickness differences. The fabrication scope for membrane phononics is therefore more closely akin to what can be achieved in the optical realm with photonic crystal waveguides, where lossless $90$ degree bends are achievable [137].
@@
For example, consider two butt-coupled waveguides of different widths, as illustrated in Fig. 29(a). CMT tells us that the coupling between the waveguides will be proportional to the overlap integral of each waveguide’s mode of oscillation. In the parallel waveguides example we could describe the perturbative field using a single travelling wave eigenmode. However, because in this case the waveguide width changes abruptly (and the eigenmodes are solutions for a fixed-width waveguide), no single eigenmode exactly represents the perturbation field. The perturbation can still be expanded into eigenmodes, but the expansion will feature infinitely many higher-order modes as shown in Fig. 29(b). The higher-order modes will propagate evanescently and do not feature in the far-field transport, but in the near-field they do play a significant role. Therefore, an analytic expression for the coupling rate between the two far-field modes would actually require summing over infinitely many overlap integrals between higher order modes [138].
@@
Figure 30: Change in resonator eigenmode from the addition of an evanescent coupler, calculated from finite element modelling with COMSOL Multiphysics [42]. (a) Fundamental mode of a 100 µm square silicon nitride resonator (b) Fundamental mode of the same resonator attached to two 30 µm wide couplers. Dashed lines highlight detail area. (c) Detail of the mode ‘leaking’ into the coupler. In this simulation the membrane is 60 nm thick, has density $3200\,\mathrm{kg\,m^{-3}}$, external stress of 1 GPa, and Young’s modulus of 230 GPa.
@@ X.1 FEM simulations and overlap integration
Our first approach is to extend our analytic expression for the coupling rates that we derived in Section VII.1, Equations (86) and (87), to the situation of a resonator coupled to a waveguide. The natural extension is to define the coupling rate of the resonator to the waveguide as:
$\sqrt{\gamma_{\mathrm{wg}}}=\frac{\Omega^{3}}{4}\iint\rho(x,y)\tilde{u}_{\mathrm{res}}\tilde{u}_{\mathrm{wg}}\,\mathrm{d}x\mathrm{d}y,$ (102)
where we normalise the amplitudes against the energy in the resonator and the power in the waveguide. For example, the normalised amplitude in the resonator $\tilde{u}_{\mathrm{res}}(x,y)$ is:
$\tilde{u}_{\mathrm{res}}(x,y)=\frac{\psi_{\mathrm{res}}(x,y)}{\sqrt{\frac{1}{2}\Omega^{2}\iint\rho(x,y)\psi_{\mathrm{res}}(x,y)^{2}\,\mathrm{d}x\mathrm{d}y}},$ (103)
where $\psi_{\mathrm{res}}(x,y)\in[-1,1]$ is the dimensionless, normalised modeshape of the resonator eigenmode, and the double integral is over all space.
@@
$\langle P_{\mathrm{wg}}\rangle=v_{g}\times\frac{1}{2}\Omega^{2}\int_{-\infty}^{\infty}\rho(x,y_{0})\psi_{\mathrm{wg}}(x,y_{0})\,\mathrm{d}x\mathrm{d}y,$ (104)
where $\psi_{\mathrm{wg}}\in[-1,1]$ is the dimensionless, normalised waveguide modeshape, and $y_{0}$ is some coordinate far enough away from the resonator that the coupling is negligible and the total power travelling through the waveguide can be evaluated with a transverse line integral.
Given the power normalisation in Eq. (104) we can write the normalised waveguide amplitude $\tilde{u}_{\mathrm{wg}}$ as:
$\tilde{u}_{\mathrm{wg}}(x,y)=\frac{\psi_{\mathrm{wg}}(x,y)}{v_{g}\times\frac{1}{2}\Omega^{2}\int_{-\infty}^{\infty}\rho(x,y_{0})\psi(x,y_{0})\,\mathrm{d}x\mathrm{d}y}.$ (105)
The numerator integral is over the coupling region, which should correspond to the area inside the tunnel barrier (if it does not, i.e. the tunnel is short enough that non-negligible acoustic fields extend through it and out the other side, we are likely outside the small-coupling regime where this theory applies).
In contrast to Section VII.1, Eq. (102) we do not subtract a reference density $\rho_{0}$. This is more consistent with suspended membranes that rely on the impedance contrast with the substrate to confine waves. Another point of contrast is that the waveguide mode is not univariate but now depends on both $x$ and $y$. This is because it breaks longitudinal symmetry where it evanescently extends into the tunnel coupler, as shown in Fig. 31.
Figure 31: Principle of FEM-aided overlap integrals. The resonator modeshape and waveguide modeshapes (here calculated for a three-port power splitter) are first independently calculated as if there was an infinite evanescent tunnel. Then the two exported modeshapes are overlaid to create a particular tunnel length, and the overlap integral is calculated.
@@ X.2 Complex eigenfrequencies
A perk of using FEM simulations is that one can avoid calculating overlap integrals entirely. This can be done by creating reflectionless boundaries on the simulation domain, for example by using a perfectly matched layer [139] or using Rayleigh damping [140, 33]. This causes energy to be lost from the model, which corresponds to the system having a complex eigenfrequency $\Omega_{r}=\sqrt{\Omega_{0}^{2}-(\gamma/2)^{2}}+i(\gamma/2)$ [24]. Here $\Omega_{0}=\sqrt{k_{\mathrm{eff}}/m_{\mathrm{eff}}}$ is the undamped eigenfrequency and $\gamma$ is the damping rate. The quality factor and damping force coefficient can be obtained from the complex eigenfrequency as [24, 42]:
$Q=\frac{\Omega_{0}}{\gamma}=\frac{\Re(\Omega_{r})}{2\Im(\Omega_{r})}.$ (106)
where $\Re(\Omega_{r})$ and $\Im(\Omega_{r})$ respectively correspond to the real and imaginary parts of the complex eigenfrequency. Eq. (106) is very useful as it allows rapid prototyping by sweeping simulation parameters and immediately observing the coupling rates without having to perform any overlap integrals.
Figure 32: Finite element modelling of a damped waveguide coupled acoustic resonator. (a) The model consists of a resonator out-coupled through two evanescent tunnels to two waveguides. Geometrical parameters such as the tunnels lengths and resonator size can be freely modified. Rayleigh damping is added to the model in the region of the out-coupling waveguides. (b) Distribution of kinetic energy density of the fundamental resonator mode, plotted on a logarithmic scale. One can observe exponential damping through the tunnels characteristic of evanescent waves, then relatively slower damping in the waveguides corresponding to the Rayleigh damping.
@@ X.3 Time domain simulations
Alternatively, instead of solving for the eigenfrequencies—a frequency-domain solution—one can simulate the decay of the resonator in the time domain, such as with a finite-difference time-domain (FDTD) code. This is the simulated version of a ringdown measurement [24, 28]. In this method the system is initialised with with nonzero displacement in the resonator and no external drive. Under these circumstances, we saw in Section IV.2 that the maximum displacement per cycle of the damped, undriven resonator $Z(t)$ will exponentially decay at a rate equal to the total coupling rate $\gamma$:s:
@@
Figure 34: The finite difference time domain method. (a) For each timestep the spatial dimensions $x$ and $y$ are discretised into a regular grid with respective grid spacings $\Delta x$ and $\Delta y$. For each combination of grid indices $i,j,n$ the solution is calculated and stored as a value $u^{n}_{ij}$. Derivatives are evaluated as linear combinations of stored solutions; for example, the red dots indicate the values that would be used to evaluate $\partial^{2}u^{n}_{ij}/\partial x^{2}$ using a second-order, centred-differencing scheme (Eq. (113)). (b) The time domain is also discretised with timestep $\Delta t$ in a mathematically equivalent way. The full domain over $t$, $x$ and $y$ can be pictured as layers of two-dimensional spatial matrices, each layer corresponding to a different timestep. Here we illustrate the $t^{\mathrm{n-1}}$, $t^{\mathrm{n}}$ and $t^{\mathrm{n+1}}$ matrices in blue, orange and green respectively.
@@
In our code we use this scheme to represent the single time derivative in the damping term. For the other derivatives we use centred difference schemes, which generally have smaller truncation errors than forward and backward schemes [141]. We substitute for the time derivative with a first order centred difference scheme:
@@
and for the spatial derivatives with a second order centred difference scheme [142]:
@@
It should be noted that Eq. (114) is not a uniquely satisfactory algorithm, and other iteration equations derived from different choices of finite difference scheme can also be used.
We enforce reflectionless boundary conditions on the simulation by employing a effective perfectly-matched layer (PML) [139], highlighted in Fig. 33(a). This is done by a gradual linear ramping up of the damping rate $\gamma$ over a distance of several wavelengths near the spatial borders. The effective PML layer emulates a spatially infinite medium by damping out all incoming acoustic energy without back-reflections.
@@
Equation (115) can be converted into an iterative formula like Eq. (114) by substituting for the additional derivatives with well-chosen finite difference schemes.

### v2

@@ IX.2 Example: Intracavity amplitude enhancement
Consider a square resonator of width $W_{\mathrm{res}}$ bookended by two waveguides of width $W_{\mathrm{wg}}$, as illustrated in Fig. 28. The resonator is impedance matched so that $\gamma_{1}=\gamma_{2}=\gamma/2$, were $\gamma_{i}$ is the coupling rate of the resonator to the $i^{\mathrm{th}}$ waveguide and $\gamma$ is the total decay rate of the resonator (we ignore intrinsic losses, i.e. $\gamma_{0}\approx 0$).
Consider the case where there is a drive in one waveguide ($a_{\mathrm{in}}=a_{\mathrm{in,1}}$ and $a_{\mathrm{in,2}}=0$), and unidirectional propagation. The drive frequency $\Omega$ is in the single mode regime of the waveguides and equals the resonator fundamental eigenfrequency ($\Delta=0$). Plugging the assumptions into Eq. (97) we find:
@@
This is equivalent to noting that the circulating power within the resonator is enhanced by a factor of $\mathcal{F}/(2\pi)$ where $\mathcal{F}$ is the resonator finesse [144].
@@
We want to know the amplitude enhancement inside the resonator, that is $Z_{0,\mathrm{res}}/Z_{\mathrm{0,\mathrm{wg}}}$. We can substitute analytic expressions for the energy in the resonator and power flow in the waveguide built into the normalisation conditions for $a$ and $a_{\mathrm{in}}$:
@@
Here $v_{g}$ is the group velocity in the waveguide from Eq. (12).
@@
In the limit where the frequency is well above cutoff, the group velocity in Eqs. (100) and (101) can be effectively approximated as $v_{g}=\sqrt{\sigma/\rho}$.
@@ IX.3 Example: Filtering
A phononic cavity can be placed into a waveguide to act as a notch pass filter. Consider a filter composed of a cavity evanescently coupled to two waveguides as in Fig. 28. Assuming it operates without reflection in the steady state ($a_{1,\mathrm{out}}=a_{2,\mathrm{in}}=\dot{a}=0$), we can solve Equations (96) and (97) to obtain $\gamma_{1}=\gamma_{0}+\gamma_{2}$. This the same impedance matching expression from Section IX.1. The total loss rate of the resonator is then $\gamma_{\mathrm{total}}=\gamma_{0}+\gamma_{1}+\gamma_{2}\simeq 2\gamma_{\mathrm{1}}$, assuming the coupling rate is engineered to be much greater than the intrinsic damping rate. This total loss rate $\gamma$ equals the full width at half maximum (in angular units) of the filter.
@@ X Numerically finding coupling rates
Numerical methods are needed because because phononic membranes can be fabricated into a large variety of shapes that resist analytic treatment. As described in Section 9, the acoustic confinement provided by the impedance mismatch between the suspended membrane and the bulk substrate is very large (from the combination of contrasting speeds of sound and from the large difference in mass between the membrane and substrate [14]) enabling low-loss acoustic wave transmission as experimentally demonstrated in our earlier work with silicon nitride membranes [12]. This low loss—much lower than what is achievable for conventional semiconductor photonic high-index waveguides [145] or phononic high-index contrast circuits[146, 5]—means that sharp bends and abrupt changes in geometry can be realised while still achieving extremely high quality factors. Indeed, a typical refractive index mismatch in silicon photonics is $n_{\mathrm{Si}}-n_{\mathrm{vacuum}}\simeq 3.5-1=2.5$, corresponding to an impedance mismatch (for a plane electromagnetic wave in a nonconductive, nonmagnetic material) of $1/2.5=0.4$. In contrast, we saw in Section I.2 that silicon nitride membranes can have more pronounced impedance mismatches with the substrate of $Z_{\mathrm{mem}}/Z_{\mathrm{sub}}\simeq 0.06$, even before considering thickness differences. The fabrication scope for membrane phononics is therefore more closely akin to what can be achieved in the optical realm with photonic crystal waveguides, where lossless $90$ degree bends are achievable [147], or in the microwave regime with hollow conductive waveguides which impose that the parallel electric field components be zero at the boundaries [148].
@@
For example, consider two butt-coupled waveguides of different widths, as illustrated in Fig. 29(a). CMT tells us that the coupling between the waveguides will be proportional to the overlap integral of each waveguide’s mode of oscillation. In the parallel waveguides example we could describe the perturbative field using a single travelling wave eigenmode. However, because in this case the waveguide width changes abruptly (and the eigenmodes are solutions for a fixed-width waveguide), no single eigenmode exactly represents the perturbation field. The perturbation can still be expanded into eigenmodes, but the expansion will feature an infinite series of higher-order modes as shown in Fig. 29(b). The higher-order modes will propagate evanescently and do not feature in the far-field transport, but in the near-field they do play a significant role. Therefore, an analytic expression for the coupling rate between the two far-field modes would actually require summing over infinitely many overlap integrals between higher order modes [149].
@@
Figure 30: Change in resonator eigenmode from the addition of an evanescent coupler, calculated from finite element modelling with COMSOL Multiphysics [42]. (a) Fundamental mode of a 100 µm square silicon nitride resonator (b) Fundamental mode of the same resonator attached to two 30 µm wide couplers. Dashed lines highlight detail area. (c) Detail of the mode ‘leaking’ into the coupler. In this simulation the membrane is 60 nm thick, has density $3200\,\mathrm{kg\cdot m^{-3}}$, external stress of 1 GPa, and Young’s modulus of 230 GPa.
@@ X.1 FEM simulations and overlap integration
Our first approach is to extend our analytic expression for the coupling rates that we derived in Section VII.1, Equations (86) and (87), to the situation of a resonator coupled to a waveguide via an in-line evanescent tunnel. Unlike Section VII.1 where we considered varying densities and modeshapes extending evanescently beyond guiding regions, here we assume a constant density $\rho$ and clamped boundary conditions, consistent with confinement by the impedance mismatch between suspended and unsuspended material. In this context, the natural extension of Eqs. (86) and (87) is to define the coupling rate of the resonator to the waveguide as:
$\sqrt{\gamma_{\mathrm{wg}}}=\frac{\rho\Omega^{3}}{4}\iint\tilde{u}_{\mathrm{res}}\tilde{u}_{\mathrm{wg}}\,\mathrm{d}x\mathrm{d}y,$ (102)
where the integral is over the coupling region, which should correspond to the area inside the tunnel barrier (if it does not, i.e. the tunnel is short enough that non-negligible acoustic fields extend through it and out the other side, we are departing from the perturbative limit upon which CMT is based). The amplitudes $\tilde{u}_{\mathrm{res}}$ and $\tilde{u}_{\mathrm{wg}}$ are normalised against the energy in the resonator and the power in the waveguide. For example, the normalised amplitude in the resonator $\tilde{u}_{\mathrm{res}}(x,y)$ is:
$\tilde{u}_{\mathrm{res}}(x,y)=\frac{\psi_{\mathrm{res}}(x,y)}{\sqrt{\frac{1}{2}\rho\Omega^{2}\iint\psi_{\mathrm{res}}(x,y)^{2}\,\mathrm{d}x\mathrm{d}y}},$ (103)
where $\psi_{\mathrm{res}}(x,y)$ is the dimensionless, normalised ($\max(|\psi_{\mathrm{res}}(x,y)|)=1$) modeshape of the resonator eigenmode, and the double integral is over the resonator area.
@@
$\langle P_{\mathrm{wg}}\rangle=v_{g}\times\frac{1}{2}\rho\Omega^{2}\int_{0}^{W_{\mathrm{wg}}}u_{\mathrm{wg}}(x,y_{0})^{2}\,\mathrm{d}x,$ (104)
where $u_{\mathrm{wg}}$ is the flexural displacement of the waveguide, $W_{\mathrm{wg}}$ is the waveguide width, and $y_{0}$ is some coordinate far enough away from the resonator that the coupling is negligible and the total power travelling through the waveguide can be evaluated with a transverse line integral. Given Eq. (104) we can write the normalised waveguide amplitude $\tilde{u}_{\mathrm{wg}}=u_{\mathrm{wg}}/\langle P_{\mathrm{wg}}\rangle$ as:
$\tilde{u}_{\mathrm{wg}}(x,y)=\frac{\psi_{\mathrm{wg}}(x,y)}{\sqrt{v_{g}\times\frac{1}{2}\rho\Omega^{2}\int_{0}^{W_{\mathrm{wg}}}\psi(x,y_{0})^{2}\,\mathrm{d}x}},$ (105)
where $\psi_{\mathrm{wg}}$ is the dimensionless, normalised ($\max(|\psi_{\mathrm{wg}}(x,y)|)=1$) waveguide modeshape.
In contrast to Section VII.1, in Eq. (102) we do not subtract a reference density $\rho_{0}$ since the density is constant everywhere. Another point of contrast is that the waveguide mode is not univariate but now depends on both $x$ and $y$. This is because it breaks longitudinal symmetry where it evanescently extends into the tunnel coupler, as shown in Fig. 31.
Figure 31: Principle of FEM-aided overlap integrals. The resonator modeshape and waveguide modeshapes (here calculated for a three-port power splitter) are first independently calculated as if there was an infinite evanescent tunnel. Then the two exported modeshapes are overlaid to create a particular tunnel length, and the overlap integral (102) is calculated.
@@ X.2 Complex eigenfrequencies
A perk of using FEM simulations is that one can avoid calculating overlap integrals entirely. This can be done by creating reflectionless boundaries on the simulation domain, for example by using a perfectly matched layer [150] or using Rayleigh damping [151, 33]. This causes energy to be lost from the model, which corresponds to the system having a complex eigenfrequency $\Omega=\sqrt{\Omega_{0}^{2}-(\gamma/2)^{2}}+i(\gamma/2)$ [24]. Here $\Omega_{0}=\sqrt{k_{\mathrm{eff}}/m_{\mathrm{eff}}}$ is the undamped eigenfrequency and $\gamma$ is the damping rate. The quality factor and damping force coefficient can be obtained from the complex eigenfrequency as [24, 42]:
$Q=\frac{\Omega_{0}}{\gamma}=\frac{\Re(\Omega)}{2\Im(\Omega)}.$ (106)
where $\Re(\Omega)$ and $\Im(\Omega)$ respectively correspond to the real and imaginary parts of the complex eigenfrequency. Eq. (106) is very useful as it allows rapid prototyping by sweeping simulation parameters and immediately observing the coupling rates without having to perform any overlap integrals.
Figure 32: Finite element modelling of a damped waveguide coupled acoustic resonator. (a) The model consists of a resonator out-coupled through two evanescent tunnels to two waveguides. Geometrical parameters such as the tunnels lengths and resonator size can be freely modified. Rayleigh damping is added to the model in the region of the out-coupling waveguides. (b) Distribution of kinetic energy density of the fundamental resonator mode, plotted on a logarithmic scale. One can observe exponential decay through the tunnels characteristic of evanescent waves, followed by damping of the excited travelling waves in the waveguides due to the Rayleigh damping.
@@ X.3 Time domain simulations
Alternatively, instead of solving for the eigenfrequencies—a frequency-domain solution—one can simulate the decay of the resonator in the time domain, such as with a finite-difference time-domain (FDTD) code. This is the simulated version of a ringdown measurement [24, 28]. In this method the system is initialised with with nonzero displacement in the resonator and no external drive. Under these circumstances, we saw in Section IV.2 that the maximum displacement per cycle of the damped, undriven resonator $Z(t)$ will exponentially decay at a rate proportional to the total coupling rate $\gamma$:
@@
Figure 34: The finite difference time domain (FDTD) method. (a) For each timestep the spatial dimensions $x$ and $y$ are discretised into a regular grid with respective grid spacings $\Delta x$ and $\Delta y$. For each combination of grid indices $i,j,n$ the solution is calculated and stored as a value $u^{n}_{ij}$. Derivatives are evaluated as linear combinations of stored solutions; for example, the red dots indicate the values that would be used to evaluate $\partial^{2}u^{n}_{ij}/\partial x^{2}$ using a second-order, centred-differencing scheme (Eq. (113)). (b) The time domain is also discretised with timestep $\Delta t$ in a mathematically equivalent way. The full domain over $t$, $x$ and $y$ can be pictured as layers of two-dimensional spatial matrices, each layer corresponding to a different timestep. Here we illustrate the $t^{\mathrm{n-1}}$, $t^{\mathrm{n}}$ and $t^{\mathrm{n+1}}$ matrices in blue, orange and green respectively.
@@
In our code we use this scheme to represent the single time derivative in the damping term. For the other derivatives we use centred difference schemes, which generally have smaller truncation errors than forward and backward schemes [152]. We substitute the time derivative with a first order centred difference scheme:
@@
and the spatial derivatives with a second order centred difference scheme [153]:
@@
It should be noted that Eq. (114) is not the only algorithm that can accurately time-evolve the system; other iteration equations derived from different choices of finite difference scheme can also be used.
We enforce reflectionless boundary conditions on the simulation by employing a effective perfectly-matched layer (PML) [150], highlighted in Fig. 33(a). This is done by a gradual linear ramping up of the damping rate $\gamma$ over a distance of several wavelengths near the spatial borders. The effective PML layer emulates a spatially infinite medium by damping out all incident acoustic energy without back-reflections.
@@
Equation (115) can be converted into an iterative formula like Eq. (114) by substituting the additional derivatives with well-chosen finite difference schemes.